OpenStudy (anonymous):

need help plzzzz will medal and fan Stan evaluated a function with a value and got 3 as the result. He wrote the entire process as a sentence: The function f(x)=x^2+2x when evaluated with 3 gives you the value of 3 or f(x)=x^2-3x when evaulated with -3 give you the value of 3 or f(x)=x^2+2x when evaluated with -3 gives you the value of 3 or f(x)=x^2-3x when evaulated with 3 gives you the value of 3

OpenStudy (anonymous):

The function f(x)=x^2+2x when evaluated with 3 gives you the value of 3

OpenStudy (anonymous):

f(x)=x^2-3x when evaulated with -3 give you the value of 3

OpenStudy (anonymous):

f(x)=x^2+2x when evaluated with -3 gives you the value of 3

OpenStudy (anonymous):

f(x)=x^2-3x when evaulated with 3 gives you the value of 3

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Luigi0210 @aaronq @hhelpplzzzz @Jesstho.-. @iGreen.